### v3

Abstract
The recent performance improvements in mixed-integer programming (MIP) went along with a significantly increased complexity of the codes of MIP solvers, which poses challenges in fixing implementation errors. In this paper, we introduce MIP-DD, a solver-independent tool, which is, to the best of our knowledge, the first open-source delta debugger for MIP. Delta debugging is a hypothesis-trial-result approach to isolate the cause of a solver failure. MIP-DD simplifies MIP instances while maintaining the undesired behavior and already supported and motivated fixes for many bugs in the SCIP releases 8.0.4, 8.1.0, and 9.0.0. This translates to an increase of approximately 71.4% more bugfixes than in the same time period before and including some fixes of long-known issues. As we highlight in selected case studies, instances triggering fundamental bugs in SCIP can typically be reduced to a few variables and constraints in less than an hour. This makes it significantly easier to manually trace and check the solution process on the resulting simplified instances. A promising future application of MIP-DD is the analysis of performance bottlenecks, which could very well benefit from simple adversarial instances.
Keywords:
Delta-debugging, mixed-integer programming1 Introduction
Over the last few decades, solvers for mixed-integer programming (MIP) have witnessed remarkable performance improvements [1], allowing them to solve even large-scale problems despite the NP-hardness of the problem class. However, these performance gains come with increasingly complex algorithms, often posing challenges in identifying and fixing implementation errors. These errors may cause the solver mistakenly claiming unboundedness or infeasibility of the problem, or even returning “optimal solutions” that are indeed infeasible or suboptimal [2, 3, 4, 5].
In this paper, we describe a new open-source tool MIP-DD [6] that can be used to facilitate and accelerate the process of locating implementation errors, commonly called “bugs”, in any complete or heuristic MIP solver. We discuss several successful case studies from the release process for Version 9.0 of the open-source MIP solver SCIP [7].
Let us first review common difficulties arising when debugging MIP solvers. Typical methods to locate bugs in the code are assertions or a debug solution mechanism. Assertions inform the user if certain essential conditions are violated at runtime. A debug solution mechanism informs the user about the invalidation of a feasible reference solution during the solving process. Nevertheless, these methods have their limitations. First, due to computational overhead, assertions and debug solution checks are usually not applied as frequently as needed to directly locate the occurrence of an error. In addition, bugs resulting from limited numerical precision might even pass assertions that incorporate certain tolerances. Second, due to the complex interaction of different solving techniques in MIP solvers, the root cause of an error may occur at an earlier stage than when an assertion fail is detected. Third, when the input instance admits multiple optimal solutions, any techniques that may cut off optimal solutions, such as strong dual reductions [8] or symmetry handling [9], might have to be turned off to guarantee that a particular debug solution is not validly excluded during the solving process, which can unintentionally work around the issue of interest.
While experienced solver developers are trained to deal with these particular difficulties, tracing back the root cause of an error can still become costly or virtually impossible due to the sheer size and runtime of input problems encountered in many bug reports. The goal of MIP-DD is to attack this size and time aspect of debugging by iteratively shrinking the problem size and the covered code base while preserving the reproducibility of the error. The reductions performed by MIP-DD then usually make it easier for a developer to isolate the actual bug on the reduced input and provide a targeted fix also for the original input.
In the literature, this approach is known as delta debugging [10], which is an automated method to isolate the cause of a software failure, driven by a hypothesis-trial-result loop. Delta debugging has been applied successfully in the SAT solver community [11, 12, 13, 4]. To the best of our knowledge, the present work describes the first application of delta debugging to mixed-integer programming and mathematical optimization in general.
The paper is organized as follows. In Section 2, we describe the structure of our delta debugger. In Section 3, we present a series of case studies to showcase the nature of different types of implementation errors and demonstrate the successful application of the delta debugger in these cases. In Section 4, we summarize different benefits of delta debugging and give an outlook on possible future improvements.
2 Structure of the Delta Debugger
A simplified workflow of MIP-DD can be seen in Figure 1. MIP-DD consists of multiple Modifiers each equipped with a unique strategy to isolate a bug. Each modifier creates a local copy of the problem and modifies it. In the general context of Delta Debugging, we then wish to test the hypothesis that the solver exhibits the same bug also on the modified problem. In order to speed up this process, modifications can be grouped in batches and tested together as one hypothesis. If the hypothesis is confirmed, i.e., the bug is reproduced, then all modifications are applied to the globally stored problem; otherwise, the modification is reverted.
In the following sections, we explain further details of the implementation. In Section 2.1, we first present the modifiers and their strategies. In Section 2.2, we discuss the API by which the modifiers interact with the solver via a solver interface. This API is able to solve the (modified) problem and needs to evaluate if a bug is reproduced. By this design choice, a broader range of bugs can be detected, see also the example discussed in Section 3.5. In Section 2.3, we motivate the ordering mechanism in which MIP-DD calls modifiers to ensure that no potential modification was missed, in particular if batches are used to speed up the process. Finally, in Section 2.4, we provide some general recommendations for using MIP-DD.
2.1 Modification Modules
MIP-DD consists of nine different modifiers each with a unique strategy to modify the problem or the settings in order to simplify the reproducing instance, reduce the code coverage, or accelerate the solving process to make the subsequent debugging easier. In order to prioritize substantial over superficial reductions, the modifiers are ordered in the following sequence:
1.
Module Constraint deletes a constraint from the problem.
2.
Module Variable fixes a variable to the value of the reference solution.
3.
Module Coefficient deletes a coefficient of a fixed variable and balances the left- and right-hand-side of the constraint .
4.
Module Fixing removes a fixed variable from the problem.
5.
Module Setting switches a parameter of the solver to a user-defined value. These values are defined in the target setting file.
6.
Module Side fixes inequalities to with being the reference solution.
7.
Module Objective sets the objective coefficient of a variable to zero.
8.
Module VarRound rounds objective coefficient and the bounds of a variable to integer values.
9.
Module ConsRound rounds the coefficients of a constraint to integer values. The right-hand and left-hand sides are adjusted accordingly.
An important requirement is that each modifier reduction must preserve the feasibility of the reference solution if it is provided. Therefore, in modifiers VarRound and ConsRound the resulting bounds and sides are relaxed to include the reference solution if necessary. Deleted variables are assumed to be fixed to the respective value of the reference solution. Optimality of the reference solution must not be preserved. Note that for infeasible problems, this approach resembles a heuristic method to identify an irreducible infeasible subset [14, 15, 16].
By default, a solve call is triggered after each single modification.
Testing each modification separately, however, can be very time-consuming.
To aggregate multiple modifications within a modifier the user can bound the number of solves per each invocation of the modifier by using the parameter nbatches
The modifiers then internally calculate how many modifications must be aggregated to a single hypothesis.
As an example consider the execution of the Constraint modifier for a MIP with constraints.
With nbatches
set to 100, the modifier aggregates 1000 constraints to one batch and tests the removal of all 1000 constraints at once, hence reducing the number of solve invocations to 100.
2.2 Solver Interface
The communication with the solver occurs via the solver interface, an API that allows easy integration of other MIP solvers. The implementation of the solver interface must report back to the modifiers whether a bug is reproduced. To do so, the solver interface should apply several checks to detect the bugs of interest. In the interface to SCIP, the following checks are applied. It is checked whether
the dual bound of the current problem cuts off the objective value of the reference solution. Under the assumption that all modifiers preserve the feasibility of the reference solution, this is a contradiction to the feasibility of the reference solution. In this case, a suboptimality bug is detected.
a solution returned by the solver is cut off by the problem restriction. In this case, an infeasibility bug is detected. For an unbounded problem, the solver sometimes additionally provides a primal ray, which is then also verified to describe an unbounded improving direction.
the primal bound is better than the evaluated objective value of the optimal solution, i.e., if the best solution is cut off.
whether an unexpected error occurs during the solving process, this is also considered a failure.
All these checks are provided by MIP-DD and can be used for any other MIP solver and can be deactivated by parameter.
To integrate a solver into MIP-DD’s API, a new class has to be created that inherits from a class SolverInterface
This solver interface interacts with the solver by
(a) parsing the settings and instance files as well as loading the data into the internal data structure (parseSettings
, readInstance
(b) translating and loading the internal settings and problem to the solver (doSetup
(c) solving the problem and checking for bugs (solve
), and
(d) writing the internal settings and problem to files (writeInstance
The solve
method returns a signed char
, where the value 0 indicates that the solver finished and no bugs are detected.
Solver internal error codes must be mapped to negative values, whereas positive values are reserved for bugs detected by the solver-independent checks provided in the abstract SolverInterface class.
MIP-DD provides template methods for checking the dual bound, the primal solutions, the primal ray, and the objective evaluation.
2.3 Workflow
When aggregating modifications in batches, valid modifications may be missed, in particular if only a small subset of the aggregated modifications is responsible for the wrong behavior of the solver. Hence, after the first invocation of a modifier, there might be valid modifications remaining in the current problem making additional calls necessary.
MIP-DD uses a stage-round mechanism to call modifiers multiple times in a different order. This mechanism is similar to the workflow of presolving, see, e.g., [17, 18]. In the -th stage only the modifiers with priority at most are called in sequence. This is repeated over several rounds until no more changes are applied. Then, the stage number is incremented or the algorithm terminates if there are no more modifiers to add to the stage cycle. Hence, if in one round the problem remains the same, the next stage is entered to include an additional modifier. Otherwise, the next round is entered to repeat the process with the same set of modifiers.
At the end of each round, the current settings-problem pair is written to files so that it can be used as an additional test instance to verify a bug fix. The user can manipulate the mechanism by defining the initial and last stage as well as the maximal number of rounds.
2.4 General Recommendations for Applying MIP-DD
Determine batch size
As explained in Section 2, the modifiers invoke the solver by default after each reduction.
Hence, a larger model size will increase the number of solve invocations.
For a more efficient process it is recommended to set parameter nbatches
as the limit of the number of solve invocations per modifier.
In this way, every successful solve invocation will simplify a substantial chunk of the problem at once.
As a rule of thumb, the longer the initial solve takes to reproduce the bug, the smaller should be the setting for parameter nbatches
, in order to decouple the expected total runtime from the original problem size.
The final instance with a possibly reduced runtime could then be further improved with a larger number of batches.
Define limits for solves
Some reductions can increase the solving time, especially when effective solving techniques are disabled. Therefore, we recommend adding time or node limits for the underlying solver both to the original and target settings. If no bug is detected within these limits, the corresponding reduction will be discarded immediately. Typically, this speeds up the process and favors the creation of easy instances.
Suppress unwanted fails
All kinds of fails are taken seriously by default.
However, it can happen that for example, a dual fail turns into a primal fail during the reduction process.
Although discovering further issues is a beneficial side effect, it is sometimes desired to avoid particular transitions of fail types in order to suppress unresolved known issues or prioritize certain bug fixes.
For this, the parameter passcodes
can be used.
All codes inside this list are additionally interpreted as correct results.
Negative values are reserved for solver-internal errors and positive values encode issues detected by MIP-DD, representing, e.g., a dual fail by 1, a primal fail by 2, and an objective fails by 3.
Exact solution
Especially for numerically sensitive issues it is important to ensure that the reference solution has feasibility violations as small as possible. To illustrate potential confusion due to an inaccurate reference solution, we consider the simple example
with optimal solutions . We assume that the solver claims the suboptimal solution to be optimal. Now, for the sake of illustration, assume that the infeasible solution would be used as a reference solution. Then might be fixed to the reference value 1, which would render the correct optimal solution. In this case, MIP-DD could interpret this correct result as a dual fail because the dual bound 0 claimed by the solver is higher than the objective value -1 of the reference solution assumed feasible.
Using the feasible solution as a reference solution instead, reliably avoids those problems. Although slight feasibility violations of the reference solution are often compensated by the problem modifications, similar issues can still occur for an almost feasible reference solution due to differing scales when applying tolerances. Therefore, using a feasible solution with minimal violation is recommended. To generate an exactly feasible solution, an inexact solution can be either polished or an exact optimal solution can be computed directly, for example by using the numerically exact version of SCIP [19, 20, 21].
3 Case Study
A preliminary version of MIP-DD has already been used to track down the majority of the bugs fixed since version 8.0.3 of the open-source solver SCIP [7]. In this section, we present selected examples to showcase different types of errors that can occur during the development of a MIP solver and demonstrate how MIP-DD supports the debugging process.
Virtually all MIP solvers employ floating-point arithmetic. In floating-point arithmetic, a limited amount of memory is allocated to represent a number. Hence, not all numbers can be represented accurately. If more digits are needed than allocated the last digits are omitted leading to a loss in precision. Among further problems, calculation in floating-point arithmetic is neither associative nor distributive meaning that is not necessarily equal to and may yield a different result than . For further information about floating-point arithmetic, we refer to [22].
To deal with these issues, floating-point MIP solvers typically define a zero tolerance and a feasibility tolerance to overcome such shortcomings. If two numbers are within -range, i.e., , they are considered equal. In SCIP, a constraint with a solution is considered feasible if
This complicates the development of a MIP solver and increases the potential for incorrect results on numerically difficult instances.
In the following case studies, we use MIP-DD 1.0 (Hash 455b7913) on a standard MacBook Air with an Apple(R) Silicon(TM) M2 2022 8‑Core CPU @ 3.49 GHz, 8‑Core GPU, and 16 GB RAM. All times are given in seconds.
Case | Original Instance | Final Instance | MIP-DD Statistics | ||||||
Vars | Conss | Nonzeroes | Vars | Conss | Nonzeroes | Rounds | MIP Solves | Time [s] | |
3.1 | 6 | 6 | 36 | 5 | 3 | 15 | 12 | 152 | 0.2 |
3.2 | 3117 | 1293 | 11751 | 7 | 4 | 11 | 66 | 69430 | 2869 |
3.3 | 653 | 745 | 1819 | 7 | 2 | 8 | 10 | 2318 | 4.1 |
3.4 | 653 | 745 | 1819 | 4 | 4 | 11 | 8 | 2198 | 4.3 |
3.5 | 204 | 104 | 940 | 0 | 0 | 0 | 5 | 638 | 3.4 |
3.1 Modifying the Settings to Detect Wrong Calculation of Activities in DualInfer
At first we want to highlight the ability of MIP-DD to identify faulty components by manipulating the settings.
For this we take a look at an instance reported by a SCIP user that has been in the backlog for over three years (SCIP git hash d4c0ad7644
Using the so-called aggressive presolve settings leads to an incorrect solution while running SCIP with default settings resulted in the correct solution.
To identify the faulty presolve routine we define the default settings as the target setting and run MIP-DD.
As a result, all settings are reset to their default except for Dual-Infer [17], which remained activated pointing to this presolver as the faulty component.
In fact, the reason for the incorrect solution is that Dual-Infer incorrectly calculated the min-max-residuum for the subset . If applied correctly, the maximal activity is calculated by
In the reduced instance
the maximum residuum is .
However, SCIP claims that the residuum is leading to incorrect assumptions in the subsequent solving process.
Since Dual-Infer applies dual reductions using the debug solution is unlikely to be effective in locating the bug even if called frequently during presolving.
The reason is an index shift where instead of and always and were applied.
This is fixed with commit cf362110
3.2 Fixing Invalid Separation of Interior Solutions
In this section we want to analyze a bug appeared on the MIPLIB [23] instance rococoC10-001000
when enabling the non-default separators intobj, closecuts, and oddcycle in the root node.
This bug can be replicated with SCIP commit 82835b0d
For this problem with minimal value 11460 a solution with value 23706 was declared optimal after 5776 LP iterations.
Applying delta debugging with disabled separators, heuristics, and presolvers as target settings, primal fails suppressed by using as passcodes
, and 1,000 batches, led to the small integer program
in less than an hour. It has the same optimal value as the original problem while a solution of value 12087 was claimed optimal already after 4 LP iterations. The only remaining non-disabling settings are presolving/maxrounds = -1, constraints/linear/maxprerounds = -1, separating/closecuts/freq = 0, and separating/intobj/freq = 0. This means that apart from some presolving only separators closecuts and intobj are required to reproduce this issue. As closecuts is known as a meta separator, these settings indicated that there might be an issue in separator intobj when called by separator closecuts [24, 25, 26]. Separator closecuts determined an interior point of the LP relaxation which is handed to intobj to be separated. Separator intobj relies on the integrality of the objective function, which is exploited by cutting off a solution with a fractional objective value. This is done by rounding this value up and using it as left-hand side of the objective row. The approach implicitly assumes that there is no feasible solution with a smaller objective value than the infeasible solution to be separated since all undiscovered better solutions would otherwise be cut off. However, for the arbitrary LP-interior points computed by separator closecuts, this does not hold in general. It turned out that the value of the separated solution is approximately 11646.03 indeed exceeding the optimal value 11460. Restricting intobj to optimal solutions of LP relaxations finally resolved this issue.
During preliminary experiments, we observed on this issue that using default settings as target settings, results in a larger final instance with 1528 variables and 301 constraints. The runtime was even comparable to the previously presented setup although disabling plugins usually leads to larger solving times of the single solves. An explanation for this behaviour might be given by the complexity of the solver under default settings which under certain circumstances requires a more complex external problem to reproduce the failing internal state. Nevertheless, targeting default settings is desirable for the generation of lightweight but flexible regular tests. And for other issues, also in this case, comparably small reproducing instances could be found.
3.3 Detecting Invalid Propagations in the Presolving Library PaPILO
It is worth mentioning that MIP-DD can also reveal further bugs that are not even reproducible initially, especially in instances that are numerically challenging in general. PaPILO [18], a presolving library for integer and linear optimization, which is part of SCIP’s presolving routines, was not called at all for the instance reported by a user. However, the delta-debugger reduced it to
where PaPILO (githash df30ae49
) states infeasibility although a solution within -tolerance exists, namely and otherwise zero.
PaPILO chooses to apply Probing [27, 17] to the variable for (LABEL:eq::3662). Probing tentatively fixes binary variables to 0 and 1, applies constraint propagation [27] and tries to conclude implication based on the outcomes of the propagation. By fixing to 0, the problem correctly propagates to infeasibility. If the variable is fixed to , propagation tries to find new bounds for the variable by resolving the constraints by :
Since the quotient exceeds the tolerance of it is floored leading to the new upper bound of -1. Since -1 is smaller than the lower bound, fixing leads to infeasibility also for . Combined with the infeasibility of Probing declares global infeasibility.
To resolve this undesired numerical effect, flooring and ceiling for integer variables are applied with more caution.
In this case, is first ceiled to 0 against .
Then it is checked if still satisfies the original constraint within .
If this is the case, this ceiled value is considered the upper bound.
Otherwise, we can safely cut off this value and the floored value is a valid upper bound.
This was fixed with the commit 019915e0
3.4 Identifying a Numerically Critical Constraint Normalization
Since PaPILO is not called in the original instance fixing the bug in PaPILO does not fix the solve status of the original instance. Therefore, re-running MIP-DD with the fixed PaPILO and suppressed primal fails results in a new reduced problem
where SCIP (git hash cf244e14
) returns a false solution.
In this problem, the second constraint leads to numerical difficulties. During presolving, all variables except are either fixed () or substituted resulting in the following problem
SCIP normalizes the remaining constraint by dividing it by leading to a right hand side that is smaller than the feasible tolerance :
Hence, (3) implies the lower bound of and can be deleted due to the redundancy allowing to choose any value for within its domain. Setting to 0 invalidates (2) since . If the normalization for vanishing left-hand sides is skipped, coefficient tightening [27, 17] for (2) is applied leading to
Immediately after the tightening, normalization is invoked again, normalizing (4) to which implies the correct fixing of to 1.
This is fixed with the commit 6b2dd64c
3.5 Fixing the Solution Count
MIP-DD can not only be used to detect primal or dual fails but can serve for testing the entire solver functionality.
For example, in the counting mode of SCIP, if a problem is already solved during presolving, a solution count of zero was claimed, even though a feasible solution was provided when using the regular optimization mode.
This bug can be reproduced with SCIP commit ca23f99f
We extended the SCIP interface to return an error if SCIP claims a solution count of zero on a feasible instance.
MIP-DD showed that this issue already occurs for the trivial problem with neither variables nor constraints due to the mishandling of a presolved problem in which all variables are fixed.
With this information the issue could be resolved with the commit 434945be
4 Conclusion and Outlook
To conclude, let us revisit different benefits of delta debugging during the development process of an MIP solver and give an outlook on possible enhancements for the future. First, as shown in Section 3, MIP-DD is able to fulfill its general purpose of reducing large instances to tiny instances while maintaining the error. This spares a lot of time during debugging by reducing the hurdles for investigating issues because all relevant calculations done by the solver on the reduced instance can usually be traced and checked manually. Further, MIP-DD can be applied even if other techniques such as providing a debug solution fail.
Second, in the context of collaborations with industry, it regularly happens that instances cannot be shared due to the General Data Protection Regulation or other confidentiality reasons. In this case, a reduced instance with renamed variable and constraints names obtained by delta debugging may barely contain any sensitive information anymore about the original instances, making the reduced instance safe for customers to share with developers.
Third, the MIP-DD itself does not detect an error, rather is the definition of an error implemented in the solver-API. By this design choice, MIP-DD can not only be applied to detect bugs caused by primal or dual infeasibility but rather to analyze any unwanted behavior specific to a certain solver. As one example, the upcoming MIP-DD 2.0 will contain, along with improvements for automatic limit settings, adaptive batch numbers, and extended arithmetical options, an interface to the numerically exact version of SCIP [28, 29, 21] and the verification software VIPR for checking validity of certificates [30]. The generation of an invalid certificate can then be chosen as another possible error to be reproduced, and reducing the instance size goes hand in hand with producing a small, invalid VIPR certificate.
In an even broader context, MIP-DD could also be utilized to track down performance bottlenecks for example by letting the solver return an error at a certain tree depth which should not be reached on a given instance. This way, issues with exploding tree sizes on instances turning out to be simple could be investigated manually, which contributes to creating more performant solvers in itself.
Acknowledgements.
The work for this article has been partly conducted within the Research Campus Modal funded by the German Federal Ministry of Education and Research (BMBF grant numbers 05M20ZBM) and the Bayreuth Research Center for Modeling and Simulation (MODUS) funded by the German Research Foundation (DFG project RA 1033/3-1).
We would like to thank all who contribute to SCIP by submitting bug tickets and allowing us to publish their instances.
Furthermore, we wish to thank Leon Eifler for helping to implement the interface to the exact version of SCIP and SoPlex, Felipe Serrano for his careful proofreading, and Armin Biere for sharing his experience and perspective on delta debugging in the SAT solver world.